Recently,theTransport Association of Zambia increased the price of bus fares for all long routes in

kwacha and 250 kwacha and clarify how bus fleetsrevenues are expected to respond to the

rises again from 250 kwacha to 300 kwacha? [5 Marks]

Zambia. Specifically, the price of tickets between Lusaka and Livingstone increased from 200 Kwacha to

c)Work out the Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200

d)Briefly discuss the Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200

between Lusaka and Livingstone.

250 kwacha. The demand equation


1
Expert's answer
2021-09-09T11:44:21-0400

"Q_d=2000-5P"


a) When P=200 kwacha, "Q_d=2000-5 \times 200=1000"

Old price quantity demanded=1000


b) When P=250 kwacha, "Q_d=2000-5 \times 250=750"

New price quantity demanded=1500


c)Price elasticity of demand

"=\frac{\frac{(Q_2-Q_1)}{(Q_1+Q_2)/2}}{\frac{(P_2-P_1)}{(P_2+P_1)/2}}"


"= \frac{\frac{(750-1000)}{(750+1000)/2}}{\frac{(250-200)}{(250+200)/2} } \\"


=-1.3


d) Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200 kwacha and 250 kwacha is elastic. Since the absolute value of elasticity is more than 1. (1.3>1).

So rise in price will lead to fall in total revenue. As rise in price will lead to more fall in quantity demanded. So total revenue will fall.